# Uniform subelliptic estimates for degenerating Fokker-Planck equations

### v1

@@ Abstract.
In this paper we expand upon recent work of Hernandez, Ranard and Riedel [5], [4], Galkowski and Zworski [3], and Li [7], by proving long time bounds for solutions to certain Fokker-Planck equations with subelliptic diffusion term. We consider the case where the jump operators $\ell_{j}$ in the Lindbladian are linear functions of $x$, and place an assumption which implies that the Hörmander condition holds for the resulting Fokker-Planck equation. By constructing a suitable parametrix for this equation we show that semiclassical derivative estimates established in [3] and [7] for elliptic diffusion also hold for subelliptic diffusion, with global bounds in $L^{p}$ for all $1\leq p\leq\infty$.
@@ 1. Introduction
Recent work by Hernandez, Ranard and Riedel [5], [4], Galkowski and Zworski [3], and Li [7], show that if a quantum system interacts suitably with an open environment, then the quantum evolution of an observable remains close to the classical evolution for times that greatly exceed the Ehrenfest time $t_{E}=\log(h^{-1})$. The observables are modeled as either semiclassical pseudodifferential operators or superpositions of Gaussian states, and the interaction with the environment by “jump” operators via the Lindblad master equation. The classical evolution is given by evolution of the symbol under the Fokker-Planck equation, which arises as the leading terms in the composition formula with the Lindbladian. We refer to the above works for a more detailed discussion of the physics, as well as numerical experiments that demonstrate the phenomenon of extended coherence.
The purpose of this note is to show that key results of these papers hold under relaxed assumptions concerning the interaction with the external environment. Roughly, in the above models the interaction involved diffusion that acted in both the space and momentum variables. The resulting Fokker-Planck equation then resembles the standard heat equation on phase space, with a small parameter in front of the diffusion that vanishes as $h\rightarrow 0$. In this work we establish the same key estimates for solutions to the Fokker-Planck equation assuming only diffusion in the momentum variables. In terms of the Lindblad jump operators, this means that the jump functions depend only on the spatial variable $x$. We assume that the diffusion spans the momentum directions, which yields that the Fokker-Planck operator is a subelliptic heat equation of Hörmander sum-of-squares type [6].
The focus of this paper is establishing the estimates of Theorem 1.1 for the Fokker-Planck equation, which we do by producing a suitable parametrix to any given order. The paremetrix is constructed via a non-isotropic heat kernel iteration, analogous to the isotropic iteration in [7]. This allows us to establish bounds in $L^{p}$ for all $1\leq p\leq\infty$. Estimates like Theorem 1.1 for $p=2$ and $p=1$ are key to the conclusions of [3] and [7], and in the final section we show how certain of their results also hold in the subelliptic case.
@@
Similar to the above papers, we consider as given a Schrödinger type Hamiltonian $p(x,\xi)=\frac{1}{2}|\xi|^{2}+V(x)$, which represents the closed quantum system. We will assume $V(x)$ is smooth and real-valued function on $\mathbb{R}^{n}$, and also assume that
@@
As in [5] we consider a family of linear jump operators, but assume that $\{\ell_{j}(x)\}_{j=1}^{J}$ depend only on the space variable $x$,
@@
If $L_{j}$ denotes multiplication by $\ell_{j}(x)$, then $L_{j},L_{j}^{*}]=0=\{\ell_{j},\bar{\ell}_{j}\}$, and the Lindbladian with coupling constant $\gamma$ simplifies to
$\mathcal{L}A=\frac{i}{h}[P,A]+\frac{\gamma}{2h}\sum_{j=1}^{J}L_{j}[A,L_{j}^{*}]+[L_{j},A]L_{j}^{*}.$
This is a model for the evolution of an observable $A$ developing under the Schrödinger system interacting stochastically with a larger environment through the collection of jump operators $L_{j}$.
The corresponding Fokker-Planck operator, which gives the leading term in $\mathcal{L}A$ when $P=p^{w}(x,hD)$ and $A=a^{w}(x,hD)$, is
@@
The main result of this paper concerns uniform regularity estimates for the solution $u=e^{tQ}u_{0}$ to the initial value problem for the Fokker-Planck equation; for $t>0$:
$\partial_{t}u=Qu,\qquad u(0,x,\xi)=u_{0}(x,\xi).$
We assume that both $\gamma,h\in(0,1]$, and let $\epsilon=\sqrt{\gamma h/2}$. If we treat $\epsilon$ and $\gamma$ as the independent variables then the estimates involve $\epsilon$, but the constants in the estimates are uniform over $\gamma,\epsilon\in(0,1]$. Li [7] considered the case that $1\leq\gamma\leq h^{-1}$, but the presence of the linear term in $\partial_{\xi}$ in (1.2), i.e.  the first sum in (1.2), leads us to require bounds on $\gamma$. For simplicity we assume $\gamma\leq 1$. If $\ell_{j}$ is real valued, however, the estimates of Theorem 1.1 hold for $\gamma\leq h^{-1}$ since in that case $\epsilon\leq 1$ and the term involving just $\gamma$ in (1.2) goes away.
Since the second order terms in $Q$ are constant coefficient, the derivative bounds on $V$ and $\ell_{j}(x)$ show that $Q,(\epsilon\partial_{x})^{\alpha}(\epsilon\partial_{\xi})^{\beta}]$ is a differential operator in $\epsilon\partial_{x}$ and $\epsilon\partial_{\xi}$ of order at most $|\alpha|+|\beta|$, with smooth bounded coefficients. The operator $e^{tQ}$ is a contraction on $L^{p}$ for all $1\leq p\leq\infty$ (see Section 6). Thus, by the Duhamel formula and Gronwall’s inequality we conclude that, for all $N$ and all $\epsilon\in[0,1]$,
(1.4) $\sum_{|\alpha|+|\beta|\leq N}\|(\epsilon\partial_{x})^{\alpha}(\epsilon\partial_{\xi})^{\beta}u(t,\cdot)\|_{L^{p}(\mathbb{R}^{n})}\\ \leq C_{N}e^{ct}\sum_{|\alpha|+|\beta|\leq N}\|(\epsilon\partial_{x})^{\alpha}(\epsilon\partial_{\xi})^{\beta}u_{0}\|_{L^{p}(\mathbb{R}^{n})}.$
Our result, analogous to ones in [3] and [7], is that these bounds hold with a constant independent of $t$, i.e.  $c=0$. The following theorem counts the order of derivatives in $x$ and $\xi$ as they are counted in the subelliptic calculus, but combining (1.4) for $t\leq 1$ with (1.6) for $T=1$ shows that (1.4) holds as written with $c=0$, for some $C_{N}$ independent of $t$.
@@
Assume that $p=\frac{1}{2}|\xi|^{2}+V(x)$ with real $V$ satisfying (1.1), and $Q$ is of the form (1.2), where (1.3) holds with non-singular $B$. Then for all $N\in\mathbb{Z}_{+}$ there is $C_{N}$ such that, with $u=e^{tQ}u_{0}$ and $\epsilon=\sqrt{\gamma h/2}$, for all $1\leq p\leq\infty$, and all $\epsilon,\gamma\in(0,1]$,
(1.5) $\sup_{0\leq t\leq\infty}\,\sum_{|\alpha|+3|\beta|\leq 3N}\|(\epsilon\partial_{\xi})^{\alpha}(\epsilon\partial_{x})^{\beta}u(t,\cdot)\|_{L^{p}(\mathbb{R}^{n})}\\ \leq C_{N}\sum_{|\alpha|+3|\beta|\leq 3N}\|(\epsilon\partial_{\xi})^{\alpha}(\epsilon\partial_{x})^{\beta}u_{0}\|_{L^{p}(\mathbb{R}^{n})}.$
Additionally, for all $N\in\mathbb{Z}_{+}$ and $T>0$ there is $C_{N,T}$ so that
(1.6) $\sup_{T\leq t\leq\infty}\,\sum_{|\alpha|+3|\beta|\leq N}\|(\epsilon\partial_{\xi})^{\alpha}(\epsilon\partial_{x})^{\beta}u(t,\cdot)\|_{L^{p}(\mathbb{R}^{n})}\leq C_{N,T}\|u_{0}\|_{L^{p}(\mathbb{R}^{n})},$
where $C_{N,T}=\mathcal{O}(T^{-\frac{1}{2}N})$ as $T\rightarrow 0^{+}$.
Our proof of Theorem 1.1 is a modification of our earlier work [9]. Unlike in that work, the Fokker-Planck equation has a distinguished time variable, which we exploit to handle the non-compact setting. Estimates are obtained through a parametrix construction, which we outline here in the case that $B$ is the identity matrix.
We start by pulling the Fokker-Planck equation back by the flow along the first-order terms in (1.2). The leading order terms in the pullback of $\partial_{t}-Q$ are then $\partial_{t}-(\partial_{\xi}+t\partial_{x})^{2}$, in the subelliptic notion of order. It is convenient to make a linear change of variables to put this in the form
@@
Care needs to be taken in that the pullback of $\partial_{\xi}$ and $\partial_{x}$ by the flow gives a frame whose second derivatives in $t$ increase unboundedly with $\xi$ due to the $\xi\cdot\partial_{x}$ term in the flow. To obtain global estimates we therefore can use only the first few derivatives in $t$, which we accomplish by using the heat kernel form of the fundamental solution rather than the pseudodifferential representation used in [9].

### v2

# Lindblad evolution with subelliptic diffusion

@@ Abstract.
We consider classical/quantum correspondence in Lindblad evolution with jump operators for which the corresponding Fokker–Planck equation is subelliptic. This allows us to consider the physical model proposed by Zurek and Paz [15], and to extend some of the recent mathematical results of Hernandez, Ranard and Riedel [9], [8], Galkowski and Zworski [5], and Li [11], where the diffusion term in the Fokker-Planck equation was assumed elliptic. We consider the case where the jump operators $\ell_{j}$ in the Lindbladian are linear functions of $x$, and place an assumption which implies that the Hörmander condition holds for the resulting Fokker-Planck equation. By constructing a suitable parametrix for this equation we show that the semiclassical derivative estimates established in [5] and [11] for elliptic diffusion also hold in the subelliptic case, with global bounds in $L^{p}$ for all $1\leq p\leq\infty$.
@@ 1. Introduction
In this paper we consider the question of the length of time for which the classical/quantum correspondence is valid for quantum systems coupled to larger environments. The basic model comes from the work of Zurek and Paz [15] and we are motivated by recent mathematical work of Hernandez, Ranard and Riedel [9], [8], Galkowski and Zworski [5], and Li [11]. The paper [9] in particular can be consulted for physical intuitions and additional indications to the literature.
To explain the setting we first consider classical evolution governed by a Hamiltonian $p=p(x,\xi)$. Here we use the mathematical notation in which $x$ and $\xi$ are respectively the position and momentum variables; a typical $p$ is given by $p(x,\xi)=\frac{1}{2}|\xi|^{2}+V(x)$, where $\frac{1}{2}|\xi|^{2}$ is the kinetic energy and $V(x)$ the potential. A classical observable (a function of position and momentum) $a=a(x,\xi)$ evolves in time according to the Hamiltonian flow $\varphi_{t}=\exp tH_{p}$, $H_{p}:=\partial_{\xi}p\cdot\partial x-\partial_{x}p\cdot\partial_{\xi}$. This means that
(1.1) $\partial_{t}a_{t}(x,\xi)=H_{p}a_{t}(x,\xi)\ \Longleftrightarrow\ a_{t}(x,\xi)=\varphi_{t}^{*}a(x,\xi):=a(\varphi_{t}(x,\xi)).$
On the classical level interaction with the environment can be modeled by adding Brownian motion in the momentum variables. This corresponds to changing the evolution equation in (1.1) to a Fokker–Planck equation:
(1.2) $\partial_{t}a_{t}(x,\xi)=H_{p}a_{t}(x,\xi)+\epsilon^{2}\Delta_{\xi}a_{t}(x,\xi).$
The connection to Brownian motion $W$ in $\xi$ comes from considering the Langevin equation
$(x^{\prime}(t),\xi^{\prime}(t))=H_{p}(x(t),\xi(t))+\epsilon\sqrt{2}\dot{W},$
so that $a_{t}(x(0),\xi(0))=\mathbb{E}(a(x(t),\xi(t)))$. See Drouot [3] and Ren–Tao [12] for recent progress on the closely related kinetic Brownian motion.
The papers [9, 5, 11] considered equations like (1.2) with $\Delta_{\xi}$ replaced by $\Delta_{x,\xi}$, that is, with comparable rates of diffusion in both $x$ and $\xi$. With this change the equation becomes a heat equation with drift, and more general cases of $p(x,\xi)$ can be handled. On the other hand, for $p=\frac{1}{2}|\xi|^{2}+V(x)$ the equation (1.2) as written leads to diffusion in $x$ through the term $\xi\cdot\partial_{x}$ in $H_{p}$, but at a different time scale than in $\xi$. Intuitively, the relation $\dot{x}=\xi$ implies that diffusion in $\xi$ at scale $\epsilon t^{1/2}$ leads to diffusion in $x$ at scale $\epsilon t^{3/2}$. This nonisotropic diffusion is captured in the propagator for (1.2) constructed in Theorem 4.2 of this paper; see (2.1) and Definition 2.1 for the definitions of $K_{0}$ and $K_{\gamma}$. It is the $\epsilon$ scaling that is key to the main results of this paper, rather than the $t$ scaling.
The condition needed on $p$ to ensure this rate of diffusion in $x$ is the step-two Hörmander condition, which is that the collection of Lie brackets $H_{p},\partial_{\xi_{j}}]$ for $1\leq j\leq n$ span the tangent space in $x$ at every point. This implies that equation (1.2) is subelliptic in the variables $(t,x,\xi)$; see Hörmander [10] and Rothschild–Stein [13]. In this paper we restrict attention to $p=\frac{1}{2}|\xi|^{2}+V(x)$ for the sake of conciseness. The spanning condition is also key to the dispersive property of the Schrödinger equation.
In the quantum framework, the Hamiltonian and the observables are Weyl quantizations (see [16, Chapter 4]) of the classical Hamiltonian and classical observables: $P=p^{\rm{w}}(x,hD)$ and $A=a^{\rm w}(x,hD)$, where $h$ is the effective Planck constant/wave length of the problem. The evolution is now governed by the Schrödinger equation:
(1.3) $\partial_{t}A_{t}=\tfrac{i}{h}[P,A_{t}]\ \Longleftrightarrow\ A_{t}=e^{itP/h}Ae^{-itP/h}.$
Classical/quantum correspondence is valid when $A_{t}$ is close to $a_{t}^{\rm{w}}(x,hD)$ from (1.1). That typically holds only for $t\leq C\log(1/h)$, the Ehrenfest time – see [16, Sect.11.4] and references given therein.
The evolved operators $A_{t}$ can be written as quantizations of classical observables, $A_{t}=a^{\rm w}(t,x,\xi,h)$. Using the calculus of Weyl operators, see [16, Theorems 4.12 and 4.18], the differential equation (1.3) can be formally written, as in [15], as
(1.4) $\begin{split}&\partial_{t}a(t,x,\xi,h)=H_{p}a(t,x,\xi,h)\\ &\ \ \ -2\sum_{k=1}^{\infty}\frac{(-1)^{k}h^{2k}}{(2k+1)!}B(D)^{2k+1}(p(x,\xi)a(t,y,\eta,h))|_{x=y,\xi=\eta},\end{split}$
where $B(D)=\tfrac{1}{2}\sigma(D_{x},D_{\xi},D_{y},D_{\eta})$ with $\sigma$ the symplectic form and $D=-i\partial$. We should stress that, mathematically, this is typically a divergent asymptotic expansion. Nevertheless, following [15] (see also [7] for an interesting numerical example in this setting), an interaction with the environment can be introduced as in (1.2) by modifying (1.4) to
(1.5) $\begin{split}&\partial_{t}a(t,x,\xi,h)=H_{p}a(t,x,\xi,h)+\varepsilon^{2}\Delta_{\xi}a(t,x,\xi,h)\\ &\ \ \ -2\sum_{k=1}^{\infty}\frac{(-1)^{k}h^{2k}}{(2k+1)!}B(D)^{2k+1}(p(x,\xi)a(t,y,\eta,h))|_{x=y,\xi=\eta}.\end{split}$
This formal equation can be presented as an evolution equation for operators $A_{t}=a^{\rm w}(t,x,hD,h)$ by noting that (see [16, (8.15)])
$(\partial_{\xi_{j}}a)^{\rm w}=-\tfrac{i}{h}[x_{j},a],$
where $x_{j}$ denotes the multiplication operator by $x_{j}$. Hence, a rigorous formulation of (1.5) reads as
$\partial_{t}A_{t}$ $=\frac{i}{h}[P,A_{t}]-\frac{\gamma}{2h}\sum_{j=1}^{n}[x_{j},[x_{j},A_{t}]]$
(1.6) $=\frac{i}{h}[P,A_{t}]+\frac{\gamma}{2h}\sum_{j=1}^{J}L_{j}[A,L_{j}^{*}]+[L_{j},A]L_{j}^{*}:=\mathcal{L}A,$
where in this case $J=n$, and
$\gamma:=2\epsilon^{2}/h,\ \ \ L_{j}:=x_{j}.$
The equation (1.6) is the general form of Lindblad evolution with a finite number of jump operators $L_{j}$. The superoperator $\mathcal{L}$ generates a semigroup which, at least formally, preserves the trace and total positivity. In fact any superoperator with these properties has to be of this form (with possibly an infinite set of jump operators) – see [9] and [5] for more on that and for references. The representation in terms of the Lindbladian is close to the point of view of Caldeira and Leggett [1], see also [6]. (We note here that as in [5] we look at the observables rather than densities.)
1.1. Assumptions on $P$ and $L_{j}$
@@
Similar to [9, 5, 11], in this paper we consider a Schrödinger type Hamiltonian $p(x,\xi)=\frac{1}{2}|\xi|^{2}+V(x)$. We will assume $V(x)$ is smooth and real-valued function on $\mathbb{R}^{n}$, and also assume that
@@
As in [9] we consider a family of jump operators $\{L_{j}\}_{j=1}^{J}$ corresponding to multiplication by a linear function $\ell_{j}$, but we assume that each $\ell_{j}$ depends only on the space variable $x$,
@@
The Lindbladian $\mathcal{L}$ is defined as in (1.6). As in the preceding discussion, this models the evolution of an observable $A$ developing under the Schrödinger system and interacting stochastically with a larger environment through the collection of jump operators $L_{j}$. We note that $\{\ell_{j},\bar{\ell}_{j}\}=0$ so the Lindblad equation has no friction; that is, the term $\mu$ in [5, (1.6)] vanishes.
Since $\{\ell_{j},\bar{\ell}_{j}\}=0$ the corresponding Fokker–Planck operator, which gives the leading term in $\mathcal{L}A$ when $P=p^{\rm w}(x,hD)$ and $A=a^{\rm w}(x,hD)$, becomes
@@
While (1.8) is a more general form than (1.2), we can treat the first sum as part of $H_{p}$ and its analysis adds no complications over that of (1.2).
1.2. Classical quantum correspondence
In Section 7 we show that certain results from [5] and [11] concerning long time classical/quantum correspondence hold for the kind of jump operators considered herein. For simplicity we focus on two particularly important cases of their results. The class of observables considered are given by semi-classical symbols based on $L^{q}$ spaces. As in those papers the point is that the agreement of Lindblad evolution and of Fokker–Planck evolutions holds for significantly longer times (e.g. $h^{-\frac{1}{2}}$ vs. $\log(1/h)$) than the agreement of the Schrödinger and Hamiltonian evolutions.
Definition 1.1.
The symbol class $S^{L^{q}}_{\rho}$ denotes symbols $a(x,\xi,h)$ such that, for all multi-indices $\alpha$, there is $C_{\alpha}$ so that
The topology on $S^{L^{q}}_{\rho}$ is that induced by the family of seminorms
The following theorem extends Theorem 2 of [11] for $\rho=\frac{1}{2}$ and $\gamma>0$ held constant, and shows that the classical and quantum evolution remain close in the trace norm for times of order $t\lesssim h^{-\frac{1}{2}}$. Here $\mathcal{L}^{1}$ denotes the trace norm for operators acting on $L^{2}(\mathbb{R}^{n})$.
@@ 1.2. Classical quantum correspondence
Assume that $a\in S^{L^{1}}_{1/2}$, and let $a(t)=e^{tQ}a$. If $A(t)$ satisfies $\partial_{t}A(t)=\mathcal{L}A(t)$ with $A(0)=\mathrm{Op}_{h}^{\rm w}(a)$, there is $C$ depending on $\|a_{0}\|_{N,S^{L^{1}}_{1/2}}$ for some $N$, such that
$\|A(t)-\mathrm{Op}_{h}^{\rm w}(a(t))\|_{\mathcal{L}^{1}}\leq Cth^{\frac{1}{2}},\quad 0\leq t<\infty.$
The paper [5] established long-time correspondence in the Hilbert-Schmidt norm $\mathcal{L}^{2}$, and yields analogous estimates when $h^{\frac{1}{3}}\lesssim\gamma\lesssim 1$. We prove the following extension of the second estimate in (6.2) of Theorem 4 in [5] to the case of the jump operators considered herein.
Theorem 1.3.
Assume that $\frac{1}{2}\leq\rho\leq\frac{2}{3}$, $a\in S^{L^{2}}_{\rho}$, and let $a(t)=e^{tQ}a$. Assume also that $\gamma=\gamma(h)$ satisfies $h^{2\rho-1}\leq\gamma\leq 1$. Then if $A(t)$ satisfies $\partial_{t}A(t)=\mathcal{L}A(t)$ with $A(0)=\mathrm{Op}_{h}^{\rm w}(a)$, there is $C$ depending on $\|a_{0}\|_{N,S^{L^{2}}_{\rho}}$ for some $N$, such that
$\|A(t)-\mathrm{Op}_{h}^{\rm w}(a(t))\|_{\mathcal{L}^{2}}\leq Cth^{2-3\rho},\quad 0\leq t<\infty.$
1.3. The subelliptic estimate
Theorem 1.2 and 1.3 are proved using the strategies of papers [11] and [5] together with global-in-time estimates for the Fokker–Planck evolution $\partial_{t}a(t)=Qa(t)$, with $Q$ given by (1.8).
We assume that both $\gamma,h\in(0,1]$, and let $\epsilon=\sqrt{\gamma h/2}$. If we treat $\epsilon$ and $\gamma$ as the independent variables then the estimates involve $\epsilon$, but the constants in the estimates are uniform over $\gamma,\epsilon\in(0,1]$. Li [11] considered the case that $1\leq\gamma\leq h^{-1}$, but the presence of the linear term in $\partial_{\xi}$ in (1.8), i.e.  the first sum in (1.8), leads us to require bounds on $\gamma$; for simplicity we assume $\gamma\leq 1$. If $\ell_{j}$ is real valued, however, the estimates of Theorem 6.1 hold for $\gamma\leq h^{-1}$, since in that case $\epsilon\leq 1$ and the first sum in (1.8) goes away.
Our main result, analogous to ones in [5] and [11], is the following global-in-time derivative estimate, valid for all $N$ and all $1\leq p\leq\infty$,
(1.10) $\sup_{0\leq t<\infty}\sum_{|\alpha|+|\beta|\leq N}\|(\epsilon\partial_{x})^{\alpha}(\epsilon\partial_{\xi})^{\beta}a(t)\|_{L^{p}(\mathbb{R}^{2n})}\\ \leq C_{N}\sum_{|\alpha|+|\beta|\leq N}\|(\epsilon\partial_{x})^{\alpha}(\epsilon\partial_{\xi})^{\beta}a(0)\|_{L^{p}(\mathbb{R}^{2n})}.$
In Section 6 we prove a more general result, Theorem 6.1, which also captures the smoothing effect of $e^{tQ}$ for $t>0$.
The proof of (1.10) is a modification of our earlier work [14] which concerned the kinetic Brownian equation as in [3]. Unlike the equations considered in that work the Fokker-Planck equation has a distinguished time variable, which we exploit to handle the non-compact setting. Additionally, rather than use a pseudodifferential construction as in [14], in this paper we construct a parametrix (to any prescribed order) via a non-isotropic heat kernel iteration, analogous to the isotropic iteration in [11]. This allows us to establish bounds in $L^{p}$ for all $1\leq p\leq\infty$.
We briefly outline the parametrix construction here in the case that $B$ is the identity matrix as in (1.2), and first consider $\epsilon=1$.
We begin by pulling the Fokker-Planck equation back by the flow along the first-order terms in (1.8). The leading order terms in the pullback of $\partial_{t}-Q$ are then $\partial_{t}-(\partial_{\xi}+t\partial_{x})^{2}$, using the subelliptic notion of order. It is convenient to make a linear change of variables to put this in the form
@@ 1.3. The subelliptic estimate
Care needs to be taken in that the pullback of $\partial_{\xi}$ and $\partial_{x}$ by the flow gives a frame whose second derivatives in $t$ increase unboundedly with $\xi$ due to the $\xi\cdot\partial_{x}$ term in the flow. To obtain global estimates we therefore can use only the first couple derivatives in $t$, which is possible since we use the heat kernel form of the fundamental solution rather than the pseudodifferential representation used in [14].
The case of $0<\epsilon\leq 1$ is handled by appropriately scaling the above parametrix construction, and taking care to note that only positive powers of $\epsilon$ arise in the various expansions.